Please help:
I have to find the center and the length of a radius of the circle:
X^2+Y^2-2x-8y+8=0
Thanks,

Answer by mukhopadhyay(490)

x^2+y^2-2x-8y+8=0
=>(x-1)^2 + (y-4)^2 = 3^2
Center is at (1,4) and radius is 3